The indicator thymol blue has two color transitions as indicated in the table at the right. Select the color expected when thymol blue is addicted to solutions having the following pH values.
aping Learning al Blue Indicator Colors The indica
Answer #1
Concepts and reason

This question deals with the determine the whether the solution is acidic or basic which is determined by calculating the pH and after we need to tell the color of pH scale as per the transition range.

Fundamentals

A solution is said to be acidic when its pH lies in the range of 0 to 6.

A solution is said to be basic when its pH lies in the range of 8 to 14.

A solution is said to be neutral when its pH is equal to 7.

0
1
2
3
4
5
6
7
Neutral
8
9
10
11
12
13
14
Acid
Base
pH Scale

The color of pH scale may vary little bit in completion, but basic indication of the color remains the same.

pH of a solution is determined by using the following formula:

pH=log[H+]{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right]

pOH is determined by using the following formula:

pOH=log[OH]{\rm{pOH}} = - \log \left[ {{\rm{O}}{{\rm{H}}^ - }} \right]

pOH determined by the above formula is then converted to pH by using the following formula:

pH+pOH=14{\rm{pH}} + {\rm{pOH}} = 14

The relation between hydronium ion [H3O+]\left[ {{{\rm{H}}_{\rm{3}}}{{\rm{O}}^{\rm{ + }}}} \right] and hydroxide ion [OH]\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] is given by the dissociation constant for pure water (Kw)\left( {{{\rm{K}}_{\rm{w}}}} \right) :

Kw=[H3O+][OH]{{\rm{K}}_{\rm{w}}} = \left[ {{{\rm{H}}_{\rm{3}}}{{\rm{O}}^{\rm{ + }}}} \right]\left[ {{\rm{O}}{{\rm{H}}^ - }} \right]

At pH = 1

The solution is completely acidic.

So, the color is Red.

At pH = 2

which is the transition value of 1.2 to 2.8.

So, there will be mixture of red and yellow which is Orange.

At pH = 6

The range is between 2.8 to 8

So, the color is and Yellow.

At pH = 10

The value is greater than 10.

So, the color should be Blue.

Ans:

At pH = 1, the color is Red.

At pH = 2, the color is Orange.

At pH = 6, the color is Yellow.

At pH = 10, the color is Blue.